OpenStudy (noreality11):

(5x+10)/(x+2)*(2x)/(4x-10) I get as far as: 5(x+2)(2x)/(x+2)2(2x-5) then I get stuck

OpenStudy (aem9296):

\[\frac{ 5x+10 }{x+2} * \frac{ 2x }{ 4x-10 }\] its this equation correct?

OpenStudy (noreality11):

yes

OpenStudy (anonymous):

okay so from the step that you got to, the (x+2) and 2 cancel. that leaves you with \[ \frac{5x }{ 2x-5 }\] are you supposed to solve for x or just simplify?

OpenStudy (noreality11):

Doesn't specify. So if x can't be solved for, i suppose just simplify.

OpenStudy (noreality11):

and thanks :)

OpenStudy (anonymous):

you're welcome, I hope it helped :)

OpenStudy (noreality11):

it helped a lot :D